# Quantum Feature Space of a Qubit Coupled to an Arbitrary Bath

### v3

@@ 3.5 Decision Tree for Noise Characterisation
We then explored the ability of simple ML algorithms to use this quantum feature space. To do this, we generated 600 different noise processes and trained a random forest with $K$-Fold cross-validation [26, 27]. For the 600 distinct noise processes, 200 processes were $1/f$, 200 were ‘$1/f$ + bump’, and 200 were coloured noise. For each noise process type, half were stationary, and half were non-stationary, i.e. 100 $1/f$ stationary noise processes and 100 $1/f$ non-stationary noise processes. We randomised several parameters between each noise profile to create distinct processes to ensure a diverse and representative dataset. Table 3 summarises the randomised parameters for each noise process and the range of values used. For the $1/f$ noise processes, the exponent ($\alpha$) was varied, and for the ‘$1/f$ + bump’, the bump location in the frequency domain ($\mu$) was varied alongside $\alpha$. For coloured noise, the division factor controlling the filter was randomised. For all non-stationary profiles, the peak position of the triangle envelope as a fraction of the total time ($T$) was randomised.
@@
To analyse the noise profiles, we used a random forest classifier from the scikit-learn library [27]. The classifier was trained on the combined vector components across all three quantum feature spaces, employing a 10-fold cross-validation approach. Each sample was labelled for stationarity and noise type, where two different trees were trained for each classification task. The decision tree classifier for stationarity achieved an average test accuracy of 0.98, and the noise type classifier achieved an average test accuracy of 0.97.
The feature importances for stationarity and noise type classification are shown in Table 4. The feature importance is calculated as the sum of the reduction in the Gini impurity index across all nodes in the tree that use the feature [27]. Interestingly, for both classification tasks, only 3-4 features were needed to explain the majority of the variance in the data, where at least one feature per observable had high importance. This demonstrates how noise process characteristics manifest themselves in the quantum feature space, which may assist in future work to clarify the mapping between a given noise process characteristic and the quantum feature space.
@@
Table 4: Feature importances represent the average contributions of the features $\alpha$, $\beta$, and $\gamma$ for each of the Quantum Feature Spaces $Q_{X}$, $Q_{Y}$, and $Q_{Z}$ to the classification task. The feature importances are normalised to sum to 1.
The high accuracy of a simple inexpensive machine learning algorithm, random forest in this case, in both tasks demonstrates the effectiveness of the quantum feature space parameters in distinguishing between different noise processes and determining their stationarity. These results further validate our approach and provide a robust method for characterising and classifying noise in quantum systems.
3.6 Properties of the Quantum Feature Space
We sought to understand the mapping between control and noise parameters and the quantum feature space. For these experiments, we explored the impact of widening control pulse widths, which simulated less ideal delta function pulses, the interpolation between noise profiles, and the impact of increasing noise energy.
We first explored the impact of widening control pulse widths on the quantum feature space. The control pulse width was varied by changing the value of $\lambda$ in Eq. 20. We set $\lambda=\frac{1}{96}$, $\frac{1}{48}$, $\frac{1}{24}$, $\frac{1}{12}$, $\frac{1}{6}$, and $\frac{1}{3}$. For these simulations, we used the ‘$1/f$ + bump’ with a bump in the PSD at the $200^{\mathrm{th}}$ bin (the unknown noise process from Sec. 3.4). Fig. 3(a) illustrates the impact of widening control pulse widths. The figure shows the interpolation between data points arising from widening control pulses is smooth and well-ordered, providing evidence that the quantum feature space is well-behaved, where it is known that smooth interpolation in a feature space is a highly desirable characteristic [28, 29, 30, 31].
@@ 3.6 Properties of the Quantum Feature Space
Results for increasing noise strengthen experiments are shown in Fig. 3(c), where increasing colour saturation represents increasing energy. For the coloured noise process and ‘$1/f$ + bump’, as the noise strength increases, all points tend toward the centre of the feature space, where $\alpha_{O}$, $\beta_{O}$, and $\gamma_{O}$ are all equal to zero. Observe that when $\alpha_{O}$, $\beta_{O}$, and $\gamma_{O}=0$, $Q_{O}D_{O}Q_{O}^{\dagger}$ becomes a zero matrix, and looking to Eq. 15, we find that the expectation of the observable is zero. This indicates that as the noise process increases in energy it saturates the system, and as such all information about the noise process and control is lost. Conversely, as the noise strength decreases, points move toward the associated identity in the quantum feature space and therefore result in the system dynamics being dominated by the control pulses.
@@
However, the ‘$1/f$ + bump’ noise process shows a more complex behaviour in Fig. 3(c) when compared to the coloured noise process. The red dots, which represent the ‘$1/f$ + bump’ noise process, rotate around $Q_{X}$ and $Q_{Y}$ as the noise strength increases. This rotation is likely due to the spectral bump in the noise profile, which introduces off-diagonal components into the noise operator. In a quantum system under control pulses, such off-diagonal elements can lead to phase shifts or rotations in the operator representation. Essentially, the result shows that the ‘$1/f$ + bump’ noise process does not simply scale the operator uniformly; it alters its orientation in the feature space.
@@
It is interesting to note that for the $1/f$ noise process, the second strongest noise point is closest to the identity (rather than the weakest noise point). One explanation of this phenomenon is the interplay between the noise’s rotational effect and its contracting effect (i.e., the tendency to “wash out” information) is nonlinear. As such, at a particular noise level, the destructive interference between the control dynamics and the noise perturbations may nearly cancel out.
@@
For both noise processes, behaviour in the $Q_{Z}$ feature space is more intuitive; points get closer to the identity as the noise decreases, indicating that the control pulses are dominating system dynamics. Overall this nuanced behaviour seen in these experiments highlights the richness of the quantum feature space and suggests that further study—perhaps through controlled variations or analytical modeling—could provide deeper insight into these dynamics.
@@
Figure 3: (a) Visualisation of the quantum feature space with widening Gaussian-shaped control pulses, with the fraction indicating the scale factor. (b) Visualisation of the quantum feature space with interpolation between noise processes implemented by utilising a linear combination of the ‘$1/f$ + bump’ and coloured Gaussian noise processes. (c) Visualisation of the quantum parameter space with two different noise processes, ‘$1/f$ + bump’ and coloured Gaussian, with colour saturation representing increasing energy.
@@ 4 Conclusion and Future Work
This work has introduced a quantum feature space based on the noise operator formalism. To create this quantum feature space, we extended the noise operator formalism to map experimental measurements directly to system-bath interaction operator parameters. We visualised the quantum feature space, mapping its behaviour to control pulse and noise process parameters. It was found that quantum feature space can be illuminating and used to identify unknown qubit noise processes without making restrictive assumptions and is a promising tool for future noise characterisation.
Future work could further refine control techniques within the quantum feature space, enhancing the ability to mitigate noise effects. In particular, exploring different types of control pulses and their parameters could provide deeper insights into optimising qubit control for various noise environments. This also includes studying a broader range of noise processes to further validate the developed methodology. Incorporating non-stationary and non-Gaussian noise models further helps us understand the quantum feature space. The methods developed could also be applied and studied in higher-dimensional quantum systems in both simulated and experimental settings.

### v4

@@ 3.6 Decision Tree for Noise Characterisation
We then explored the ability of simple ML algorithms to use this QFS. To do this, we generated 600 different noise processes and trained a random forest with $K$-Fold cross-validation [james2013introduction, scikit-learn]. For the 600 distinct noise processes, 200 processes were $1/f$, 200 were ‘$1/f$ + bump’, and 200 were coloured noise. For each noise process type, half were stationary, and half were non-stationary, i.e. 100 $1/f$ stationary noise processes and 100 $1/f$ non-stationary noise processes. We randomised several parameters between each noise profile to create distinct processes to ensure a diverse and representative dataset. Table 5 summarises the randomised parameters for each noise process and the range of values used. For the $1/f$ noise processes, the exponent ($\alpha$) was varied, and for the ‘$1/f$ + bump’, the bump location in the frequency domain ($\mu$) was varied alongside $\alpha$. For coloured noise, the division factor controlling the filter was randomised. For all non-stationary profiles, the peak position of the triangle envelope as a fraction of the total time ($T$) was randomised.
@@
To analyse the noise profiles, we used a random forest classifier from the scikit-learn library [scikit-learn]. The classifier was trained on the combined vector components across all three QFSs, employing a 10-fold cross-validation approach. Each sample was labelled for stationarity and noise type, where two different trees were trained for each classification task. The decision tree classifier for stationarity achieved an average test accuracy of 0.98, and the noise type classifier achieved an average test accuracy of 0.97.
The feature importances for stationarity and noise type classification are shown in Table 6. The feature importance is calculated as the sum of the reduction in the Gini impurity index across all nodes in the tree that use the feature [scikit-learn]. Interestingly, for both classification tasks, only 3-4 features were needed to explain the majority of the variance in the data, where at least one feature per observable had high importance. This demonstrates how noise process characteristics manifest themselves in the QFS, which may assist in future work to clarify the mapping between a given noise process characteristic and the QFS.
@@
Table 6: Feature importances represent the average contributions of the features $\alpha$, $\beta$, and $\gamma$ for each of the QFSs $\tilde{X}$, $\tilde{Y}$, and $\tilde{Z}$ to the classification task. The feature importances are normalised to sum to 1.
For the sake of completeness, we also trained a K-Nearest Neighbors [peterson2009k] and Logistic regression algorithms [hosmer2013applied]. The results are shown in Table 8. As can be seen there is high accuracy across all the simple and inexpensive ML algorithms in both tasks which demonstrates the effectiveness of the QFS parameters in distinguishing between different noise processes and determining their stationarity. This is, as the feature space is well-defined and captures the essential characteristics of the noise, simple algorithms can easily leverage this structure and achieve high classification accuracy. These results further validate our approach and provides more robust methods for characterising and classifying noise in quantum systems.
Table 7: Algorithm Accuracy Comparison for Classification Tasks
Algorithm Stationarity Classification Noise Type Classification
Decision Tree 0.98 0.97
Logistic Regression 0.84 0.91
K-Nearest Neighbors 0.96 0.94
Table 8: Comparison of classification accuracies for different algorithms. The decision tree classifier achieved the highest accuracy in both tasks, stationarity and noise type classification, with K-Nearest Neighbours following closely. Logistic regression performed well but was less accurate than the other two algorithms.
3.7 Properties of the QFS
We sought to understand the mapping between control and noise parameters and the QFS. For these experiments, we explored the impact of widening control pulse widths, which simulated less ideal delta function pulses, the interpolation between noise profiles, and the impact of increasing noise energy.
We first explored the impact of widening control pulse widths on the QFS. The control pulse width was varied by changing the value of $\lambda$ in Eq. 22. We set $\lambda=\frac{1}{96}$, $\frac{1}{48}$, $\frac{1}{24}$, $\frac{1}{12}$, $\frac{1}{6}$, and $\frac{1}{3}$. For these simulations, we used the ‘$1/f$ + bump’ with a bump in the PSD at the $200^{\mathrm{th}}$ bin (the unknown noise process from Sec. 3.4). Fig. 3(a) illustrates the impact of widening control pulse widths. The figure shows that the interpolation between data points arising from widening control pulses is smooth and well-ordered, providing evidence that the QFS is well-behaved, where it is known that smooth interpolation in a feature space is a highly desirable characteristic [bengio2013representation, radford2015unsupervised, guo2024smooth, higgins2017beta].
@@ 3.7 Properties of the QFS
Results for increasing noise strengthen experiments are shown in Fig. 3(c), where increasing colour saturation represents increasing energy. For the coloured noise process and ‘$1/f$ + bump’, as the noise strength increases, all points tend toward the centre of the feature space, where $\alpha_{O}$, $\beta_{O}$, and $\gamma_{O}$ are all equal to zero. Observe that when $\alpha_{O}$, $\beta_{O}$, and $\gamma_{O}=0$, $\tilde{O}$ becomes a zero matrix, and looking to Eq. 13, we find that the expectation of the observable is zero. This indicates that as the noise process increases in energy, it saturates the system, and as such, all information about the noise process and control is lost. Conversely, as the noise strength decreases, points move toward the associated identity in the QFS and therefore result in the system dynamics being dominated by the control pulses.
@@
However, the ‘$1/f$ + bump’ noise process shows a more complex behaviour in Fig. 3(c) when compared to the coloured noise process. The red dots, which represent the ‘$1/f$ + bump’ noise process, rotate around $\tilde{X}$ and $\tilde{Y}$ as the noise strength increases. This rotation is likely due to the spectral bump in the noise profile, which introduces off-diagonal components into the noise operator. In a quantum system under control pulses, such off-diagonal elements can lead to phase shifts or rotations in the operator representation. Essentially, the result shows that the ‘$1/f$ + bump’ noise process does not simply scale the operator uniformly; it alters its orientation in the feature space.
@@
It is interesting to note that for the $1/f$ noise process, the second strongest noise point is closest to the identity (rather than the weakest noise point). One explanation of this phenomenon is that the interplay between the noise’s rotational effect and its contracting effect (i.e., the tendency to “wash out” information) is nonlinear. As such, at a particular noise level, the destructive interference between the control dynamics and the noise perturbations may nearly cancel out.
@@
For both noise processes, behaviour in the $\tilde{Z}$ feature space is more intuitive; points get closer to the identity as the noise decreases, indicating that the control pulses are dominating system dynamics. Overall, this nuanced behaviour seen in these experiments highlights the richness of the QFS and suggests that further study—perhaps through controlled variations or analytical modelling—could provide deeper insight into these dynamics.
@@
Figure 3: (a) Visualisation of the QFS with widening Gaussian-shaped control pulses, with the fraction indicating the scale factor. (b) Visualisation of the QFS with interpolation between noise processes implemented by utilising a linear combination of the ‘$1/f$ + bump’ and coloured Gaussian noise processes. (c) Visualisation of the quantum parameter space with two different noise processes, ‘$1/f$ + bump’ and coloured Gaussian, with colour saturation representing increasing energy.
@@ 4 Conclusion and Future Work
This work has introduced a QFS based on the noise operator formalism. To create this QFS, we extended the noise operator formalism to map experimental measurements directly to system-bath interaction operator parameters. We visualised the QFS, mapping its behaviour to control pulse and noise process parameters. It was found that QFS can be illuminating and used to identify unknown qubit noise processes without making restrictive assumptions, and is a promising tool for future noise characterisation.
Future work could further refine control techniques within the QFS, enhancing the ability to mitigate noise effects. In particular, exploring different types of control pulses and their parameters could provide deeper insights into optimising qubit control for various noise environments. This also includes studying a broader range of noise processes to further validate the developed methodology. Incorporating non-stationary and non-Gaussian noise models further helps us understand the QFS. The methods developed could also be applied and studied in higher-dimensional quantum systems (i.e. multi-qubit systems) in both simulated and experimental settings, to investigate if the QFS can be used for higher-dimensional systems.